# On Geometric Evolution and Microlocal Regularity of the Navier–Stokes Equations

### v2

@@ Abstract
We develop a geometric and microlocal framework for the Navier–Stokes equations by lifting the dynamics to the cosphere bundle of a Riemannian manifold. In this formulation, the velocity field and vorticity are represented as microlocal distributions whose evolution is governed by a linear transport–dissipation system generated by a canonical dynamical vector field. We introduce microlocal amplitudes, directional energy functionals, and monotone volume invariants on the compact phase space, which quantify directional concentration and alignment mechanisms associated with potential loss of regularity. The viscous term induces an effective geometric diffusion on the cosphere bundle, yielding closed differential inequalities in a geometric setting. To capture the interaction between fluid deformation and geometry, we define an effective connection and curvature tensor encoding the influence of the symmetric velocity gradient. This structure gives rise to a Ricci–type microlocal geometric evolution that constrains directional stretching and excludes extreme angular concentration compatible with viscous dissipation. While the present results do not resolve the global regularity problem, they provide a coherent geometric mechanism that severely restricts admissible blow–up scenarios, reformulating the regularity question as a problem of dissipative stability on a compact phase space.
@@
Navier–Stokes equations, microlocal analysis, cosphere bundle, geometric regularity criteria, effective Ricci curvature, dissipative dynamical systems, differential forms.
@@ I Introduction: The Regularity Problem and Geometric Flows
The question about the existence and uniqueness of the solution space for the Navier-Stokes equations is one of the most significant and open research topics in mathematics and theoretical physics [11, 40], with a vast range of applications. Geometric and covariant formulations on Riemannian manifolds provide a coherent framework independent of coordinates [7, 13]. This clearly explains differential operators, the energy structure, the role of curvature, and the symmetries of the system through Killing fields [6]. This is because Hamiltonian formulations and differential forms, using weighted and covariant tensors with exterior derivatives [30], ensure the method works naturally, making diffusion and boundary terms appear more clearly in components with non-zero curvature.
@@
In this work, we reformulate the incompressible Navier–Stokes equations as a dissipative dynamical system defined globally on the cosphere bundle $S^{*}M$. Our goal is to classify blow-up mechanisms and obtain geometric criteria for conditional regularity using variational and Fourier methods. In this way, we find inequalities and invariants to show how the physical solution is reconstructed from the dynamics in $S^{*}M$.
Microlocal lifting approach
The microlocal lifting to the cosphere bundle $S^{*}M$ adopted in this paper should be understood as a parametric refinement and compactification of the dynamics on the base manifold $M$. All geometric objects—including the effective connection $\nabla^{u}$, its torsion and curvature, and the metric $g^{\prime}$—are defined on $M$, with coefficients depending smoothly on $(x,\xi)\in S^{*}M$. The cotangent variable $\xi$ serves to encode directional and spectral information, but no geometric structure is imposed on the tangent bundle of $S^{*}M$ itself. This viewpoint allows us to combine microlocal analysis with intrinsic geometric evolution without introducing additional bundle-level connections.
@@ III Microlocal Lifting and the Cosphere Bundle
$\nabla^{u}g^{\prime}=-\bigl(A(u)\cdot g^{\prime}+g^{\prime}\cdot A(u)\bigr)$ (41)
@@ IV Geometric Evolution Operators and Microlocal Representation of Solutions
By homogeneity, the field $X_{u}(x,\xi)$ is tangent to $S^{*}M$ and therefore defines a well-defined dynamics on the cosphere bundle, where $\|\xi\|_{g^{\prime}}$ is preserved. This will be useful because, at this point, it is beneficial to perform a Fourier transform on the system [5, 6]. Since the classical Fourier transform is not globally defined on compact manifolds, we employ a microlocal transform defined locally in normal coordinates and extended globally as a function on $T^{*}M$. For a 1-form $\phi\in\Omega^{1}(M)$, we define
@@
where $\chi$ is a smooth cutoff function. Restricting by homogeneity to $S^{*}M$, we obtain $\tilde{\phi}\in C^{\infty}(S^{*}M)$. The transform (68) should be understood as a directional microlocal projection rather than a pointwise invertible transform, because of that, the horizontal lift is uniquely determined once the effective connection is fixed, although it is not canonical with respect to the background metric.
Applying the microlocal transform to the incompressible covariant system and projecting onto transverse directions to eliminate the pressure term, we obtain the microlocal equation
@@
Equation (69) is a linear non-autonomous dissipative system on $S^{*}M$. Here, the term corresponding to the pressure $dP$ does not contribute to the system’s dynamics, acting as a Lagrange multiplier that imposes the incompressibility condition, existing only in the radial direction of the cotangent bundle. Therefore, we define the incompressible microlocal projector
@@
which yields the conditions
$\Pi_{\xi}(\xi)=0,\quad\Pi_{\xi}(dP)=0$ (72)
This causes the pressure term to vanish when restricting the equation to the cosphere bundle and projecting onto transverse directions because it acts on the base manifold variables and does not generate transverse components in phase space. In fact, the condition of the divergence transform also justifies the defined projection and confirms that the dynamics evolve naturally in $S^{*}M$.
@@
In order to solve the microlocal system on the cosphere bundle $S^{*}M$, we define the Green operator $G(t,s)$ as the fundamental solution of the non-autonomous dissipative equation [25, 38]. Let $Q(t)=\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}$ be the differential operator acting on sections of the bundle. The Green operator satisfies
$\left(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}\right)G(t,s)=\delta(t-s)\mathbb{I}$ (73)
@@
where $\mathbb{I}$ is the identity operator on $C^{\infty}(S^{*}M)$.
@@
Let $\Psi_{t,s}:S^{*}M\longmapsto S^{*}M$ be the flow generated by the dynamical vector field $X_{u}$. This flow represents the lifting of the fluid’s trajectories to the phase space. For any initial microlocal distribution $\tilde{\phi}_{0}$, the pure transport part without dissipation is given by the pull-back $(\Psi_{t,s}^{*}\tilde{\phi}_{0})(x,\xi)=\tilde{\phi}_{0}(\Psi_{s,t}(x,\xi))$.
@@
In the general case, let $\tilde{f}\in C^{\infty}(S^{*}M\times N)$ represent the total microlocal source term, encompassing both external forces and non-linear residuals with N{N} is a general manifold representing additional degrees of freedom, parameters or internal symmetries of the system. In the relativistic limit, $N$ is identified with the timelike component of the worldline, and the physical fields are recovered by the push-forward integrating over the fibers. Then, the Green operator is constructed by coupling the geometric transport with the exponential dissipation factor induced by the viscosity and the metric.
$G(t,s)[\tilde{f}](x,\xi)=\exp\left(-\int_{s}^{t}\nu\|\xi(\tau)\|_{g^{\prime}}^{2}\,d\tau\right)\tilde{f}(\Psi_{s,t}(x,\xi))$ (74)
@@
  • $\xi(\tau)$ denotes the evolution of the covector along the integral curves of $X_{u}$ in the cotangent bundle.
  • $\tilde{f}$ is understood as a section whose values in N{N} are preserved along the flow $\Psi_{s,t}$ unless acted upon by internal dynamics, being part of the phase space for the system, describing the Fourier transform of the external forces and non-linear residuals of curvature.
To verify that $G(t,s)$ is the fundamental solution, we observe, at $t=s$, the flow $\Psi_{s,s}$ is the identity and the integral in the exponent vanishes, thus $G(s,s)=\mathbb{I}$. Applying the operator $\partial_{t}+\mathcal{L}_{X_{u}}$ to the transported term, we have by the property of the Lie derivative and the flow
$(\partial_{t}+\mathcal{L}_{X_{u}})\left[\tilde{f}(\Psi_{s,t}(x,\xi))\right]=0$ (75)
which follows from the fact that $\Psi_{s,t}$ is the solution to the transport equation associated with $X_{u}$. Differentiating the exponential factor with respect to $t$ yields the term $\nu\|\xi\|_{g}^{2}$. Summing these contributions, we obtain
$(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2})G(t,s)\\ =(-\nu\|\xi\|_{g^{\prime}}^{2}+\nu\|\xi\|_{g^{\prime}}^{2}+\mathcal{L}_{X_{u}}-\mathcal{L}_{X_{u}})G(t,s)=0$ (76)
for $t>s$ and it converges for the delta distribution at $t=s$. The operator thus acts as a low-pass filter in the phase space, where the dissipation follows the geometric trajectories.
Thus, the solution to the transformed microlocal system can be formally written by solving for $\tilde{\phi}$ over the time variable as
$\tilde{\phi}(t)=\mathcal{T}\bigg[\exp\left(-\int_{t_{0}}^{t}\big(\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}\big)\,ds\right)\tilde{\phi}_{0}\bigg]\\ +\int_{t_{0}}^{t}G(t,s)\big[\tilde{R}(s)+\tilde{F}(s)\big]\,ds$ (77)

### v3

@@ Abstract
We develop a geometric and microlocal framework for the Navier–Stokes equations by lifting the dynamics to the cosphere bundle of a Riemannian manifold. In this formulation, the velocity field and vorticity are represented as microlocal distributions whose evolution is governed by a linear, non-autonomous transport–dissipation system generated by a canonical dynamical vector field on the compact phase space $S^{*}M$. We introduce microlocal amplitude functionals, a directional entropy functional, and monotone geometric invariants that quantify angular concentration and alignment mechanisms associated with potential loss of regularity. The viscous term induces an effective geometric diffusion on the cosphere bundle, yielding closed differential inequalities expressed in intrinsic geometric form. To capture the interaction between fluid deformation and geometry, we define an effective connection and an associated curvature tensor encoding the influence of the symmetric velocity gradient, giving rise to a Ricci–type microlocal geometric evolution that constrains directional stretching. The framework establishes a necessary and sufficient geometric equivalence criterion: finite-time singularity occurs if and only if at least one of three microlocal controls fails—deformation integrability, entropy boundedness, or lifted energy boundedness. The symmetry structure reveals a dimensional reduction mechanism: as effective fiber dimension increases, the cosphere bundle undergoes a symmetry lock, forcing microlocal distributions toward isotropy and topologically obstructing the formation of angular singularities. While these results do not resolve the global regularity problem, they reformulate it as a question of dissipative stability and spectral control on a compact, symmetry-constrained microlocal dynamical system, severely restricting admissible blow–up scenarios through geometric and representation-theoretic constraints.
@@
Navier–Stokes equations, microlocal analysis, cosphere bundle, geometric regularity criteria, effective Ricci curvature, dissipative dynamical systems, entropy functionals, symmetry lock mechanism.
@@ I Introduction: The Regularity Problem and Geometric Flows
The question about the existence and uniqueness of the solution space for the Navier-Stokes equations is one of the most significant and open research topics in mathematics and theoretical physics [11, 40], with a vast range of applications. Geometric and covariant formulations on Riemannian manifolds provide a coherent framework independent of coordinates [7, 13]. This formulation provides a natural geometric interpretation of differential operators, the energy structure, curvature effects, and symmetries through Killing fields. [6]. This is due to the fact that Hamiltonian formulations and differential forms, using weighted and covariant tensors with exterior derivatives [30], ensure the method works naturally, making diffusion and boundary terms appear more clearly in components with non-zero curvature.
@@
In this work, we reformulate the incompressible Navier–Stokes equations as a dissipative dynamical system defined globally on the cosphere bundle $S^{*}M$. This formulation allows us to encode both spatial and directional information within a unified geometric framework.
Our objective is to investigate the structure of potential blow–up mechanisms through microlocal and geometric methods, and to derive conditions under which directional concentration can be controlled. In particular, we introduce geometric functionals and differential inequalities that relate deformation, transport, and dissipation at the microlocal level.
These constructions suggest a geometric equivalence principle in which finite–time singularity is linked to the breakdown of microlocal stability and deformation integrability. While the present work does not resolve the global regularity problem, it provides a coherent framework that restricts admissible singular behaviors and clarifies the geometric mechanisms underlying regularity.
In what follows, we use $|\cdot|$ to denote pointwise norms or fiberwise magnitudes, and $\|\cdot\|$ to denote functional norms on the relevant spaces.
@@ III Microlocal Lifting and the Cosphere Bundle
$\nabla^{u}g^{\prime}=-\bigl[A(u)g^{\prime}+g^{\prime}A(u)\bigr]$ (41)
@@ IV Geometric Evolution Operators and Microlocal Representation of Solutions
By homogeneity, the field $X_{u}(x,\xi)$ is tangent to $S^{*}M$ and therefore defines a well-defined dynamics on the cosphere bundle, where $\|\xi\|_{g^{\prime}}$ is preserved. This property motivates the introduction of a microlocal transform adapted to the compact setting. Since the classical Fourier transform is not globally defined on compact manifolds, we employ a microlocal transform defined locally in normal coordinates and extended globally as a function on $T^{*}M$ [5, 6].
IV.1 Microlocal Fourier Transform and Tempered Distributions
To place the microlocal lift in a fully controlled functional framework, we work in the space of tempered distributions. Let $M$ be a smooth $n$-dimensional manifold. In local charts, we define the Schwartz space
$\mathcal{S}(\mathbb{R}^{n})=\Big\{f\in C^{\infty}(\mathbb{R}^{n}):\sup_{x\in\mathbb{R}^{n}}(1+|x|)^{N}|\partial^{\alpha}f(x)|<\infty,$ (68)
$\forall N\in\mathbb{N},\alpha\in\mathbb{N}_{0}^{n}\Big\}$
and its dual $\mathcal{S}^{\prime}(\mathbb{R}^{n})$ of tempered distributions. Globally, the space $\mathcal{S}^{\prime}(M)$ is defined as the set of distributions whose local restrictions are tempered:
$\mathcal{S}^{\prime}(M)=\left\{u\in\mathcal{D}^{\prime}(M)\mid u\circ\varphi_{k}^{-1}\in\mathcal{S}^{\prime}(\mathbb{R}^{n})\forall k\in\{0,\dots,l\}\right\}$ (69)
where the definition is independent of the particular choice of the atlas $\mathcal{A}$.
For a 1-form $\phi\in\Omega^{1}(M)$ with sufficient regularity, we define the microlocal lift by
@@ IV.1 Microlocal Fourier Transform and Tempered Distributions
where $\chi$ is a smooth cutoff function supported in a normal neighborhood of the origin in $T_{x}M$, and $\langle\xi,y\rangle$ denotes the natural pairing between covectors and vectors. The exponential map $\exp_{x}:T_{x}M\longmapsto M$ is well-defined in a neighborhood of zero, and the cutoff ensures that the integral is taken over a region where normal coordinates are valid.
Remark 1 (Interpretation of the microlocal transform).
The transform (70) should be understood as a directional microlocal projection rather than a pointwise invertible transform. It extracts the Fourier content of $\phi$ along the direction $\xi$ in the cotangent space. The horizontal lift of $\phi$ to $S^{*}M$ is uniquely determined once the effective connection is fixed, although it is not canonical with respect to the background metric alone.
The microlocal Fourier transform in local coordinates is given by
$\mathcal{F}_{x}u(\xi)=\int_{\mathbb{R}^{n}}e^{-ix\cdot\xi}u(x)\,dx$ (71)
which extends continuously as a map
$\mathcal{F}:\mathcal{S}^{\prime}(\mathbb{R}^{n})\longmapsto\mathcal{S}^{\prime}(\mathbb{R}^{n})$ (72)
For $u\in H^{s}(\mathbb{R}^{n})$ with $s\in\mathbb{R}$, the Sobolev norm is characterized via the Fourier transform:
$\|u\|_{H^{s}}^{2}=\int_{\mathbb{R}^{n}}(1+|\xi|^{2})^{s}|\widehat{u}(\xi)|^{2}\,d\xi$ (73)
This shows that the microlocal lift inherits Sobolev control through fiberwise polynomial weights in the frequency variable $\xi$.
Proposition 1 (Microlocal Sobolev inequality).
For any $s\geq 0$ and $u\in H^{s}(M)$, the lifted object $\tilde{u}$ satisfies
$\|\tilde{u}\|_{L^{2}(S^{*}M)}\leq\sigma_{s}\|u\|_{H^{s}(M)}$ (74)
where $\sigma_{s}$ depends only on $s$ and the geometry of $M$.
Proof.
In local coordinates, the microlocal lift coincides with the restriction of the Fourier transform to the unit sphere in frequency space. By the trace theorem for Fourier multipliers and the coarea formula, restriction to $\|\xi\|_{g}=1$ lowers the Sobolev order by $1/2$. The global bound follows by summing over a finite coordinate cover using a partition of unity, noting that coordinate changes on $M$ induce bounded transformations on Sobolev spaces. ∎
This inequality ensures that microlocal regularity is controlled by classical Sobolev norms on the base manifold.
IV.2 Microlocal Projection and Incompressibility
Applying the microlocal transform to the incompressible covariant system and projecting onto transverse directions eliminates the pressure term. We obtain the microlocal equation
@@ IV.2 Microlocal Projection and Incompressibility
Equation (75) is a linear, non-autonomous, dissipative system on $S^{*}M$. The term corresponding to the pressure $dP$ does not contribute to the dynamics when restricted to the cosphere bundle, as it acts only in the radial direction of the cotangent bundle and is eliminated by the transversality condition.
To formalize this, we define the incompressible microlocal projector
@@
This projector satisfies the key properties:
$\Pi_{\xi}(\xi)=0,\qquad\Pi_{\xi}(dP)=0$ (78)
The first property reflects the fact that $\xi$ is radial in the fiber, while the second shows that the pressure gradient, being a pure differential, has no transverse component on the cosphere bundle. Consequently, when we restrict to $S^{*}M$ and project onto $\xi^{\perp}$, the pressure term vanishes identically, and the incompressibility condition transforms into the orthogonality constraint $\langle\tilde{\phi},\xi\rangle_{g^{\prime}}=0$.
IV.3 Estimates for the Commutator $d,\mathcal{L}_{X_{u}}+\nu\|\xi\|^{2}_{g^{\prime}}]$
A key technical ingredient in the microlocal analysis is the control of the commutator between the exterior derivative $d$ and the lifted evolution operator. This commutator encodes the deviation from exact compatibility between differentiation and microlocal transport.
We define the commutator operator
$R_{u}:=[d,\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}]=[d,\mathcal{L}_{X_{u}}]+[d,\nu\|\xi\|_{g^{\prime}}^{2}]$ (79)
For the first term, we use the identity for the commutator of the exterior derivative with the Lie derivative:
$d,\mathcal{L}_{X_{u}}]=\mathcal{L}_{dX_{u}}$ (80)
which follows from Cartan’s formula $\mathcal{L}_{X}=d\iota_{X}+\iota_{X}d$ and the fact that $d^{2}=0$.
For the second term, the exterior derivative of the scalar function $\nu\|\xi\|_{g^{\prime}}^{2}$ produces a 1-form, and the commutator with multiplication acts as a wedge product:
$d,\nu\|\xi\|_{g^{\prime}}^{2}]\cdot(\cdot)=d(\nu\|\xi\|_{g^{\prime}}^{2})\wedge(\cdot)$ (81)
Combining these, we obtain
$R_{u}=\mathcal{L}_{dX_{u}}+d(\nu\|\xi\|_{g^{\prime}}^{2})\wedge(\cdot)$ (82)
To estimate the operator norm of $R_{u}$ on Sobolev spaces, we note that:
  • The term $\mathcal{L}_{dX_{u}}$ is a first-order differential operator whose coefficients depend on $\nabla X_{u}$, and hence on $\nabla u$ and $\nabla^{2}u$ through the definition of the lifted field $X_{u}$.
  • The term $d(\nu\|\xi\|_{g^{\prime}}^{2})$ involves derivatives of the effective metric $g^{\prime}$ with respect to the base variables, controlled by $\|\nabla u\|_{L^{\infty}}$ due to the velocity-dependence of $g^{\prime}$.
Proposition 2 (Commutator estimate).
For any $\tilde{\phi}\in H^{1}(S^{*}M)$, the commutator $R_{u}$ satisfies the estimate
$\|R_{u}\tilde{\phi}\|_{L^{2}(S^{*}M)}\leq\sigma_{0}\left(\|\nabla u\|_{L^{\infty}}+\|\nabla^{2}u\|_{L^{\infty}}\right)\|\tilde{\phi}\|_{H^{1/2}(S^{*}M)}\\ +\nu\sigma_{1}\|\tilde{\phi}\|_{H^{1/2}(S^{*}M)}$ (83)
where $\sigma_{0},\sigma_{1}$ depend only on the background geometry.
Proof.
By symbolic calculus on compact manifolds, $R_{u}$ is a zeroth-order pseudodifferential operator with principal symbol controlled by $\nabla u$ and $\nabla^{2}u$. Standard symbol estimates yield
$\|R_{u}\|_{L^{2}\to L^{2}}\leq\sigma_{0}(\|\nabla u\|_{L^{\infty}}+\|\nabla^{2}u\|_{L^{\infty}}+\nu)$ (84)
For the refined estimate involving $H^{1/2}$ on the right-hand side, we use interpolation between $L^{2}$ and $H^{1}$, noting that $R_{u}$ loses at most half a derivative due to its structure as a commutator of first-order operators. ∎
This loss of half a derivative is absorbed by the dissipative part of the evolution operator in the energy estimates. More precisely, for any $\varepsilon>0$, we can establish the absorption inequality:
Lemma 1 (Absorption of commutator terms).
For any $\varepsilon>0$ and $\tilde{\phi}\in H^{1}(S^{*}M)$, there holds
$|\langle R_{u}\tilde{\phi},\tilde{\phi}\rangle_{L^{2}(S^{*}M)}|\leq\varepsilon\nu\|\tilde{\phi}\|_{H^{1}(S^{*}M)}^{2}\\ +\sigma_{\varepsilon}\left(1+\|\nabla u\|_{L^{\infty}}^{2}\right)\|\tilde{\phi}\|_{L^{2}(S^{*}M)}^{2}$ (85)
where $C_{\varepsilon}$ depends on $\varepsilon$ and the background geometry.
Proof.
By Proposition 2 and Young’s inequality,
[…]
$\leq\sigma_{0}(\|\nabla u\|_{L^{\infty}}+\|\nabla^{2}u\|_{L^{\infty}}+\nu)\|\tilde{\phi}\|_{H^{1/2}}\|\tilde{\phi}\|_{L^{2}}$ (87)
By interpolation, $\|\tilde{\phi}\|_{H^{1/2}}\leq\|\tilde{\phi}\|_{H^{1}}^{1/2}\|\tilde{\phi}\|_{L^{2}}^{1/2}$. Applying Young’s inequality with exponent $\varepsilon$:
$\|\tilde{\phi}\|_{H^{1/2}}\|\tilde{\phi}\|_{L^{2}}$ $\leq\|\tilde{\phi}\|_{H^{1}}^{1/2}\|\tilde{\phi}\|_{L^{2}}^{3/2}$ (88)
$\leq\frac{\varepsilon}{2\sigma_{0}\nu}\|\tilde{\phi}\|_{H^{1}}^{2}+\frac{\sigma_{0}}{2\varepsilon\nu}\|\tilde{\phi}\|_{L^{2}}^{2}$ (89)
Absorbing the first term into the dissipation yields the result. ∎
This ensures that the commutator $R_{u}$ acts as a lower-order perturbation in the energy estimates and does not interfere with the dissipative control provided by the viscous term $\nu\|\xi\|_{g^{\prime}}^{2}$.
IV.4 Green Operator and Solution Representation
@@
To solve the microlocal system on the cosphere bundle $S^{*}M$, we construct the Green operator $G(t,s)$ as the fundamental solution of the non-autonomous dissipative equation. Let
$Q(t)=\mathcal{L}_{X_{u}(t)}+\nu\|\xi\|_{g^{\prime}}^{2}$ (90)
denote the differential operator acting on sections of the bundle. The Green operator satisfies the defining equation
$\left(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|^{2}_{g^{\prime}}\right)G(t,s)=\delta(t-s)I$ (91)
@@
where $I$ is the identity operator on $C^{\infty}(S^{*}M)$ and $\delta(t-s)$ is the Dirac distribution in time.
Let $\Psi_{t,s}:S^{*}M\to S^{*}M$ denote the flow generated by the dynamical vector field $X_{u}$, defined by
$\frac{d}{dt}\Psi_{t,s}(x,\xi)=X_{u}(\Psi_{t,s}(x,\xi)),\qquad\Psi_{s,s}(x,\xi)=(x,\xi)$ (92)
@@
This flow represents the lifting of the fluid’s trajectories to the phase space $S^{*}M$. For any initial microlocal distribution $\tilde{\phi}_{0}$, the pure transport part (without dissipation) is given by the pullback
$(\Psi_{t,s}^{*}\tilde{\phi}_{0})(x,\xi)=\tilde{\phi}_{0}(\Psi_{s,t}(x,\xi))$ (93)
where $\Psi_{s,t}=\Psi_{t,s}^{-1}$ is the backward flow.
@@
In the general case, let $\tilde{f}\in C^{\infty}(S^{*}M\times[t_{0},T])$ represent the total microlocal source term, encompassing both external forces and non-linear residuals with curvature corrections. The Green operator is constructed by coupling the geometric transport with the exponential dissipation factor induced by viscosity:
$G(t,s)[\tilde{f}](x,\xi)=\exp\left(-\int_{s}^{t}\nu\|\xi(\tau)\|^{2}_{g^{\prime}}\,d\tau\right)\tilde{f}(\Psi_{s,t}(x,\xi))$ (94)
@@
  • $\xi(\tau)$ denotes the evolution of the covector along the integral curves of $X_{u}$ in the cotangent bundle,
  • The integral in the exponent accumulates the dissipation along the microlocal trajectory.
Proposition 3 (Verification of the Green operator).
The operator $G(t,s)$ defined in (94) satisfies equation (91).
Proof.
At $t=s$, the flow $\Psi_{s,s}$ is the identity and the integral in the exponent vanishes, thus $G(s,s)=I$.
For $t>s$, we verify that $G(t,s)$ solves the homogeneous equation
$\left(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|^{2}_{g^{\prime}}\right)G(t,s)[\tilde{f}]=0$ (95)
Applying $\partial_{t}+\mathcal{L}_{X_{u}}$ to the transported term $\tilde{f}(\Psi_{s,t}(x,\xi))$:
By the property of the Lie derivative and the flow, we have
$(\partial_{t}+\mathcal{L}_{X_{u}})[\tilde{f}(\Psi_{s,t}(x,\xi))]=0$ (96)
since $\Psi_{s,t}$ is the solution to the transport equation associated with $X_{u}$.
Differentiating the exponential factor with respect to $t$:
$\frac{d}{dt}\exp\left(-\int_{s}^{t}\nu\|\xi(\tau)\|^{2}_{g^{\prime}}\,d\tau\right)\\ =-\nu\|\xi(t)\|^{2}_{g^{\prime}}\exp\left(-\int_{s}^{t}\nu\|\xi(\tau)\|^{2}_{g^{\prime}}\,d\tau\right)$ (97)
Summing these contributions:
$(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|^{2}_{g^{\prime}})G(t,s)[\tilde{f}]\\ =\left(-\nu\|\xi\|^{2}_{g^{\prime}}+\nu\|\xi\|^{2}_{g^{\prime}}\right)G(t,s)[\tilde{f}]=0$ (98)
At $t=s$, the exponential factor equals 1 and the flow is the identity, so $G(s,s)=I$, confirming the delta-function normalization. ∎
The operator thus acts as a low-pass filter in phase space: dissipation accumulates along geometric trajectories, with high-frequency (large $\|\xi\|_{g^{\prime}}$) components being exponentially suppressed.
IV.5 Solution Formula and Reconstruction
The solution to the microlocal system can be formally written using the Green operator and time-ordering:
$\tilde{\phi}(t)=\mathcal{T}\left\{\exp\left[-\int_{t_{0}}^{t}\left(\mathcal{L}_{X_{u}(s)}+\nu\|\xi\|^{2}_{g^{\prime}}\right)ds\right]\right\}\tilde{\phi}_{0}\\ +\int_{t_{0}}^{t}G(t,s)\left[\tilde{R}(s)+\tilde{F}(s)\right]ds$ (99)